# Linear and Sublinear Diversities

### v2

@@ 3 Embedding into linear and sublinear diversities
We now turn our attention from linear and sublinear diversities to the questions of when finite diversities can be isometrically embedded within linear or sublinear diversities. Questions about embedding of metric spaces have, of course, been central to metric geometry and its applications, particularly after Linial et al. [18] demonstrated the link between approximate embeddings and combinatorial optimization algorithms on graphs. We showed in [7] that an analogous link holds between approximate embeddings of diversities and combinatorial optimization algorithms on hypergraphs. Here we only consider embeddings without distortion, that is, exact rather than approximate embeddings.
A map $f:X_{1}\mapsto X_{2}$ between two diversities $(X_{1},\delta_{1})$ and $(X_{2},\delta_{2})$ is an isometric embedding if $\delta_{2}(f(A))=\delta_{1}(A)$ for all finite $A\subseteq X_{1}$. We say that a finite diversity $(X,\delta)$ is linear-embeddable if there is an isometric embedding from $(X,\delta)$ to a linear diversity on $\mathbb{R}^{k}$ for some $k$ and sublinear-embeddable if there is an isometric embedding to some sublinear diversity on $\mathbb{R}^{k}$, for some $k$. At this stage we allow the dimension $k$ to be arbitrary.
Theorem 11 gives a characterization of linear-embeddability while Theorem 12 gives a characterization of sublinear-embeddability. Minkowski diversities and negative type diversities were reviewed earlier. We first establish a lemma on finite diversities that are embeddable in extremal linear diversities.
Lemma 10.
If $(\mathbb{R}^{k},\delta)$ is an extremal linear semidiversity, and $X\subseteq\mathbb{R}_{k}$ is finite then $(X,\delta)$ is Minkowski embeddable with a simplex kernel.
@@
By Theorem 7 $(\mathbb{R}^{k},\delta)$ is the Minkowski semidiversity with kernel $K=\mathrm{conv}(W)+H^{\perp}$, where $W$ is a set of affinely independent vectors lying in a subspace $H$. Let $T$ be an orthogonal matrix so that $TH=\mathrm{span}(\{e_{1},\ldots,e_{j}\})$ and $TH^{\perp}=\mathrm{span}(\{e_{j+1},\ldots,e_{j}\})$. Let $T_{H}$ be the first $j$ rows of $T$, so that $T_{H}H=\mathbb{R}^{j}$, $T_{H}H^{\perp}=\{0\}$ and $T_{H}K=\mathrm{conv}(T_{H}W)$ is a full-dimensional simplex in $\mathbb{R}^{j}$. Then for all $\lambda$, $A+x\subset\lambda K$ for some $x\in\mathbb{R}^{k}$ if and only if $T_{H}A+y\subset\lambda\mathrm{conv}(T_{H}W)$ for some $y\in\mathbb{R}^{j}$. So $\delta(A)=\delta_{\mathrm{conv}(T_{H}W)}(T_{H}A)$ for all finite $A$, as required. ∎
@@
Let $(X,\delta)$ be a finite diversity. The following are equivalent:
@@
    $(X,\delta)$ has negative type.
@@
(i)$\Rightarrow$(ii) Without loss of generality assume $X\subseteq\mathbb{R}^{k}$ where $(\mathbb{R}^{k},\delta)$ is a linear diversity. From Theorem 8 we have that any linear diversity can be expressed as a convex combination of extremal linear semidiversities. By Lemma 10 each of these extremal linear semidiversities can be expressed as a Minkowski diversity with a simplex, each of which has negative type by Theorem 17 in [3]. As the set of negative type diversities forms a convex cone, $(X,\delta)$ also has negative type.
(ii)$\Leftrightarrow$(iii) This is Theorem 17 in [3].
(iii)$\Rightarrow$(i) Theorem 8 shows that any Minkowski diversity with a simplex kernel (being a trivial example of a weighted average of such diversities) is linear. Therefore, if $(X,\delta)$ is embeddable in a Minkowski diversity with a simplex kernel, it also embeddable in a linear diversity. ∎
@@
Let $(X,\delta)$ be a finite diversity. The following are equivalent:
@@
  3. (iii)
    $(X,\delta)$ is the maximum of a collection of negative type diversities.
@@
(ii)$\Rightarrow$(i) If $(X,\delta)$ is embeddable as a Minkowski diversity, then it is sublinear-embeddable, since by Theorem 2.4 of [3] all Minkowski diversities are sublinear.
(i)$\Rightarrow$(ii) Let $(X,\delta)$ be a sublinear-embeddable. We may assume $X$ is a subset of $\mathbb{R}^{k}$ where $(\mathbb{R}^{k},\delta)$ is a sublinear diversity. By Theorem 9 there is a family of linear semidiversities $\delta_{\gamma}$ for $\gamma\in\Gamma$ such that $\delta(A)=\max\delta_{\gamma}(A)$. Since $X$ is finite, it has a finite number of subsets, and so we may assume that $\Gamma$ is finite. By Proposition 4.1 (a) in [3] if two finite diversities are Minkowski embeddable, then so is their maximum, and hence the same is true of any finite number of finite Minkowski embeddable diversities. Therefore $(X,\delta)$ is Minkowski embeddable.
(i)$\Rightarrow$(iii) We may assume $X\subseteq\mathbb{R}^{k}$ where $(\mathbb{R}^{k},\delta)$ is a sublinear diversity. By Theorem 9 there is a family of linear semidiversities $\delta_{\gamma}$ for $\gamma\in\Gamma$ such that $\delta(A)=\max\delta_{\gamma}(A)$. Since $X$ is finite, it has a finite number of subsets, and so we may assume that $\Gamma$ is finite. By Theorem 11, each of $(X,\delta_{\gamma})$ is negative type, and therefore $(X,\delta)$ is the maximum of a collection of negative type diversities.
(iii)$\Rightarrow$(ii) Suppose $(X,\delta)$ is the maximum of a collection of negative type diversities. By Theorem 11 $(X,\delta)$ can then be represented as the maximum of a collection of Minkowski diversities, and since $X$ is finite, we may assume the collection is finite. By Proposition 4.1 (a) in [3] the maximum of a finite collection of Minkowski embeddable diversities is Minkowski embeddable. ∎

### v3

@@ 3 Embedding into linear and sublinear diversities
We now turn our attention from linear and sublinear semidiversities to characterizations of when finite semidiversities can be embedded into linear or sublinear diversities. Questions about embedding of metric spaces have, of course, been central to metric geometry and its applications, particularly after Linial et al. LinialLondonEtal95 demonstrated the link between approximate embeddings and combinatorial optimization algorithms on graphs. We showed in BryantTupper14 that an analogous link holds between approximate embeddings of diversities and combinatorial optimization algorithms on hypergraphs. This has been used by jozefiak2023diversity to obtain the best current approximation bounds for sparsest cut on some classes on hypergraphs. Here we only consider embeddings without distortion, that is, exact rather than approximate embeddings.
A map $f:X_{1}\to X_{2}$ between two semidiversities $(X_{1},\delta_{1})$ and $(X_{2},\delta_{2})$ is an embedding if $\delta_{2}(f(A))=\delta_{1}(A)$ for all finite $A\subseteq X_{1}$. Since $f$ is not required to be injective, a semidiversity can be embedded into a diversity. We say that a finite semidiversity $(X,\delta)$ is linear-embeddable if there is an embedding from $(X,\delta)$ to a linear diversity on $\mathbb{R}^{k}$ for some $k$ and sublinear-embeddable if there is an embedding to some sublinear diversity on $\mathbb{R}^{k}$, for some $k$. At this stage we allow the dimension $k$ to be arbitrary.
Theorem 9 gives a characterization of linear-embeddability while Theorem 10 gives a characterization of sublinear-embeddability. Minkowski diversities and diversities of negative type were reviewed earlier.
Before we state our characterization of linear-embeddable semidiversities we need to revisit diversities of negative type, extending the results of WuBryantEtal19 to semidiversities. Recall that a finite semidiversity has negative type if it satisfies $\sum_{A,B}x_{A}x_{B}\,\delta(A\cup B)\leq 0$ for all vectors $x$ indexed by the subsets of $X$ with $x_{\emptyset}=0$ and $\sum_{A}x_{A}=0$.
Proposition 8.
Let $(X,\delta)$ be a finite semidiversity. Then $(X,\delta)$ has negative type if and only if there is an embedding of $(X,\delta)$ into a finite diversity of negative type.
@@
[…]
$\sum_{A,B\subseteq X}x_{A}x_{B}\delta(A\cup B)$ $=\sum_{I,J\subseteq Y}\sum_{A:\phi(A)=I}\sum_{B:\phi(B)=J}x_{A}x_{B}\,\widetilde{\delta}(\phi(A\cup B))$
$=\sum_{I,J\subseteq Y}\left(\sum_{A:\phi(A)=I}x_{A}\right)\left(\sum_{B:\phi(B)=J}x_{B}\right)\widetilde{\delta}(I\cup J)$
$=\sum_{I,J\subseteq X}\tilde{x}_{I}\tilde{x}_{J}\widetilde{\delta}(I\cup J)$
$\leq 0,$
[…]
Let $y$ be a zero-sum vector, indexed by subsets of $Y$, such that $y_{\emptyset}=0$. Define a vector $x$, indexed by subsets of $X$, where
$x_{A}=\begin{cases}y_{A}&A\subseteq Y\\ 0&\mbox{otherwise.}\end{cases}$
Then $x$ has zero sum, $x_{\emptyset}=0$ and
$\sum_{A,B\subseteq Y}y_{A}y_{B}\widetilde{\delta}(A\cup B)=\sum_{A,B\subseteq X}x_{A}x_{B}\delta(A\cup B)\leq 0.$
[…]
Suppose that $A\subseteq X$. Then
$\delta(A)\leq\delta(A\cup\phi(A))\leq\delta(\phi(A))+\sum_{a\in A}\delta(\{a,\phi(a)\})=\delta(\phi(A))=\widetilde{\delta}(\phi(A))$
and
$\widetilde{\delta}(\phi(A))=\delta(\phi(A))\leq\delta(\phi(A)\cup A)\leq\delta(A)+\sum_{a\in A}\delta(\{a,\phi(a)\})=\delta(A),$
[…]
@@
Let $(X,\delta)$ be a finite semidiversity. The following are equivalent:
@@
    $(X,\delta)$ is of negative type.
@@
[…]
$\delta_{\mbox{\footnotesize{neg}}}(A)=\sum_{i=1}^{\ell}\max_{a\in A}a_{i}-\min_{a\in A}\left\{\sum_{i=1}^{\ell}a_{i}\right\}.$
We will determine an embedding $\phi\colon X\rightarrow\mathbb{R}^{\ell}$ such that
$\delta(A)=\sum_{i=1}^{\ell}\max_{a\in A}\phi(a)_{i}\ \ \ \ \text{and}\ \ \ \ \ \ \sum_{i=1}^{\ell}\phi_{i}(x)=0,$
thereby showing that $(X,\delta)$ is embeddable in a diversity of negative type, and is therefore of negative type by Proposition 8.
By Theorem 5, $\delta$ is given by
$\delta(A)=\int_{\mathbb{S}^{k-1}}h_{A}(u)\,\mathrm{d}\nu(u)$
for a positive measure $\nu$ on the sphere satisfying $\int_{\mathbb{S}^{k-1}}u\,\mathrm{d}\nu(u)=0$. Every $u\in\mathbb{S}^{k-1}$ induces an ordering $\sigma\colon\{1,\ldots,n\}\rightarrow X$ of the points in $X$ so that
$u^{T}x_{\sigma(1)}\leq\cdots\leq u^{T}x_{\sigma(n)}.$ (9)
We denote the order for a given $u$ by $\sigma_{u}$ (breaking ties using the lexicographic order). Let the set of all $u$ such that $\sigma_{u}=\sigma$ be denoted $S_{\sigma}$. If we integrate (9) over $S_{\sigma}$ with respect to $\nu$ we obtain
$u_{\sigma}^{T}x_{\sigma(1)}\leq\cdots\leq u_{\sigma}^{T}x_{\sigma(n)}$ (10)
where $u_{\sigma}=\int_{S_{\sigma}}u\,d\nu(u)$. So for a given set $A$, and for all $u\in S_{\sigma}$, the $a\in A$ that maximizes $u^{T}a$ is the same as the $a$ that maximizes $u_{\sigma}^{T}a$. Let us denote this value $a$ by $m_{A,\sigma}$.
For any set $A\subseteq X$ we have
$\delta(A)=\int_{\mathbb{S}^{k-1}}h_{A}(u)d\nu(u)=\sum_{\sigma}\int_{S_{\sigma}}h_{A}(u)d\nu(u)=\sum_{\sigma}\int_{S_{\sigma}}u^{T}m_{A,\sigma}\,d\nu(u)=\sum_{\sigma}u_{\sigma}^{T}m_{A,\sigma}=\sum_{\sigma}\max_{a\in A}u_{\sigma}^{T}a.$
We define a map $\phi\colon X\rightarrow\mathbb{R}^{\Sigma}$ where $\Sigma$ is the set of all permutations of members of $X$. We define $\phi(a)_{\sigma}=u_{\sigma}^{T}a$ for $a\in X$. So $\delta(A)=\sum_{\sigma}\max_{a\in A}\phi(a)_{\sigma}$
Now we observe that for all $a\in X$
$\sum_{\sigma}\phi(a)_{\sigma}=\sum_{\sigma}u_{\sigma}^{T}a=\sum_{\sigma}\int_{S_{\sigma}}u^{T}ad\nu(u)=\left(\int_{\mathbb{S}^{k-1}}ud\nu\right)^{T}a=0^{T}a=0.$
So $(X,\delta)$ is embeddable in $(\mathbb{R}^{|\Sigma|},\delta_{\mbox{\footnotesize{neg}}})$ which is a diversity of negative type, and therefore $(X,\delta)$ is a semidiversity of negative type.
(ii) $\Rightarrow$(iii). By Theorem 5.2 of bryant2023diversities , any diversity of negative type can be embedded into a Minkowski diversity with a simplex kernel. By Proposition 8 this also holds for semidiversities of negative type.
(iii) $\Rightarrow$ (i) By Theorem 5.2 of bryant2023diversities , a Minkowski diversity with simplex kernel is linear, so $(X,\delta)$ is linear-embeddable. ∎
@@
Let $(X,\delta)$ be a finite semidiversity. The following are equivalent:
@@
    There is a finite collection $\{(X,\delta_{\gamma}):\gamma\in\Gamma\}$ of semidiversities of negative type such that
    $\delta(A)=\max\{\delta_{\gamma}:\gamma\in\Gamma\}$
    for all $A\subseteq X$.
  3. (iii)
@@
[…]
$\delta(A)=\max\{\delta_{\gamma}(A):\gamma\in\Gamma\}$
can also be embedded in a Minkowski diversity.
[…]